# Quantum Machine Learning and Grover’s Algorithm for Quantum Optimization of Robotic Manipulators

### v1

@@ III-B2 Step 2: Quantum Machine Learning for Kinematic Approximation
where $\hat{\bm{P}}$ is the position observable corresponding to $\bm{p}(\bm{q},\bm{d})$ in the classical formulation, and $U_{\text{QML}}(\bm{\theta})$ is a parameterized quantum circuit that entangles qubits to capture complex kinematic relationships, similar to neural network layers in robotic learning tasks. The quantum model is trained by minimizing the discrepancy with the true FK:
@@
where $\bm{p}_{k}^{\text{QML}}$ is the QML-predicted position for configuration $\lvert k\rangle$,and $f_{\text{FK}}(\bm{q}^{(k)},\bm{d}^{(k)})\|^{2}$ is the true FK. This QML optimization [33] uses gradient descent on classical computers to tune $\bm{\theta}$, while quantum hardware evaluates circuit outputs—similar to training a robotic simulator to match real-world kinematics, but leveraging quantum parallelism. For classification in QML, refer to [29].
III-B3 Step 3: Robotic Cost Hamiltonian Formulation
@@
Here, the Hamiltonian $\hat{H}$ acts like a potential energy landscape in robotics, where lower eigenvalues correspond to better configurations—quantum measurement samples from this landscape. Alternative observables include orientation error $\hat{O}_{\mathrm{rot}}=\sum_{k}d_{R}(\bm{R}_{k},\bm{R}_{\mathrm{tgt}})^{2}\,\lvert k\rangle\langle k\rvert$, manipulability $\hat{O}_{\mathrm{manip}}=-\sum_{k}\sqrt{\det(\bm{J}_{k}\bm{J}_{k}^{\top})}\,\lvert k\rangle\langle k\rvert$, and effort/energy $\hat{O}_{\mathrm{energy}}=\sum_{k}\|\bm{\tau}_{k}\|_{2}^{2}\,\lvert k\rangle\langle k\rvert$, as summarized in Table I. The cost Hamiltonian is:
$\hat{H}_{\text{cost}}=w\cdot\hat{O}_{\text{task}}$ (8)
where $\hat{O}_{\text{task}}$ is the primary observable, and $w$ is a weight. For pose optimization:
$\hat{H}_{\text{pose}}=\alpha_{p}\cdot\hat{O}_{\text{pose}}$ (9)
consistent with the classical weight $\alpha_{p}$ from Section III-A. The expectation value $\langle\psi\rvert\hat{H}\lvert\psi\rangle$ gives the average cost over superposed configurations, analogous to Monte Carlo sampling in robotic optimization.
III-B4 Step 4: Grover-Optimized Configuration Search
Grover’s algorithm amplifies quantum states corresponding to configurations satisfying the task objective. For IK problems, this is:
$\|f_{\text{FK}}(\bm{q},\bm{d})-\bm{p}_{\text{target}}\|^{2}\leq\epsilon$ (10)
where $\epsilon$ is the position tolerance. Grover provides quadratic speedup for finding optimal configurations among many possibilities, similar to prioritized sampling in robotic reinforcement learning. Algorithm 1 implements this search:
Input: Target pose $T_{\text{target}}$, tolerance $\epsilon$, max iterations $K_{\max}$
Output: Optimal configuration $(\bm{q}^{*},\bm{d}^{*})$
Initialize uniform superposition: $\lvert\psi_{0}\rangle=H^{\otimes N}\lvert 0\rangle^{\otimes N}$;
Train $U_{\text{QML}}$ to approximate forward kinematics;
Construct $\hat{H}_{\text{cost}}$ for target objective (Step 3);
Define IK oracle: $\mathcal{O}_{\text{IK}}\lvert k\rangle=\begin{cases}-\lvert k\rangle&\|f_{\text{QML}}(\lvert k\rangle)-\bm{p}_{\text{target}}\|^{2}\leq\epsilon\\ \lvert k\rangle&\text{otherwise}\end{cases}$;
Define diffusion operator: $D=2\lvert\psi_{0}\rangle\langle\psi_{0}\rvert-I$;
for $k\leftarrow 1$ to $\min\left(\left\lfloor\frac{\pi}{4}\sqrt{2^{N}}\right\rfloor,K_{\max}\right)$ do
   Apply $\mathcal{O}_{\text{IK}}$ (phase flip for solutions);
   Apply $D$ (probability amplification);
end for
Measure final state $\rightarrow$ obtain $b^{*}$;
Decode $b^{*}\rightarrow(\bm{q}^{*},\bm{d}^{*})$;
return $(\bm{q}^{*},\bm{d}^{*})$;
Algorithm 1 Grover-Enhanced Configuration Search
@@ III-B4 Step 4: Grover-Optimized Configuration Search
where $\alpha_{p},\alpha_{R}>0$ balance translational and rotational units and $\epsilon$ is the admissible pose error. Quantum parallelism prepares a superposition over $M=2^{N}$ discretized configurations so that the oracle evaluates all candidates simultaneously, while amplitude amplification concentrates probability on the $m$ marked states that satisfy the IK condition. The number of Grover iterations is chosen near the optimum $K\approx\bigl{\lfloor}(\pi/4)\sqrt{M/m}\,\bigr{\rfloor}$, yielding quadratic speedup over classical search. For general optimization, the oracle employs the cost Hamiltonian:
@@
III-B5 Step 5: Classical Verification and Feasibility Assessment
@@
$e_{\text{actual}}=\bigl{\|}\,f_{\text{FK}}^{\text{analytical}}(\bm{q}^{*},\bm{d}^{*})-T_{\text{target}}\,\bigr{\|}_{F}.$ (13)
In the general case, the oracle marks low-cost states through the cost Hamiltonian, subsuming the IK predicate. After computing $e_{\text{actual}}$, feasibility is confirmed by checking hard constraints (joint limits, closures, collisions) and task tolerance. Failing candidates prompt adjustment of $\epsilon$, discretization, or QML training before reapplying Grover. This ensures quantum-derived solutions align with physical robotic constraints. To demonstrate implementation and effectiveness, we present three case studies translating robotic optimization problems into quantum pipelines. The examples progress from a 1-DoF planar arm to a 2-DoF serial manipulator, and to a dual-arm grasping scenario using two 2-DoF manipulators. These are chosen for conceptual clarity, presenting each methodology stage in simple form, building on the prior framework.
@@ IV Case Studies
IV-A Case Study 1: 1-DoF Manipulator for Pose Optimization
We begin with a planar single-revolute-joint manipulator, formulated as a discrete search for compatibility with quantum methods and consistency with higher-dimensional optimizations. The objective is to optimize the link length $l_{1}$ and joint angle $\theta_{1}$ to achieve the target end-effector position $\bm{p}_{\text{target}}=(p_{x},p_{y})$ within tolerance $\epsilon$. Although optimizing $l_{1}$ is straightforward here, it establishes a scalable foundation for hyperparameter optimization. The quantum pipeline executes the five stages: 1) Discretization & Encoding: parameter spaces are discretized with $l_{1}\in[l_{\min},l_{\max}]$ encoded via $n_{l}$ qubits and $\theta_{1}\in[0,2\pi)$ via $n_{\theta}$ qubits, yielding $N=n_{l}+n_{\theta}$ qubits and $M=2^{N}$ configurations. 2) QML Kinematic Approximation: parameterized circuit $U_{\text{QML}}(\bm{\theta})$ learns the mapping: $\bm{p}(l_{1},\theta_{1})=[l_{1}\cos\theta_{1},\ l_{1}\sin\theta_{1}]^{\top}$ with observable $\hat{\bm{P}}$ extracting positional data. 3) Cost Hamiltonian Formulation: the position-error observable implements the objective: $\hat{H}_{\text{cost}}=\alpha_{p}\sum_{k=0}^{M-1}\|\bm{p}_{k}^{\text{QML}}-\bm{p}_{\text{target}}\|^{2}\lvert k\rangle\langle k\rvert$ with $\alpha_{R}=0$ for this degenerate orientation case. 4) Grover-Optimized Search: Oracle $\mathcal{O}_{\text{IK}}$ marks valid configurations: $\mathcal{O}_{\text{IK}}\lvert k\rangle=-\lvert k\rangle\quad\text{iff}\quad\|\bm{p}_{k}^{\text{QML}}-\bm{p}_{\text{target}}\|\leq\epsilon$ with $\sim\lfloor(\pi/4)\sqrt{M/m}\rfloor$ iterations for $m$ solutions. 5) Classical Verification: Decoded solution $(l_{1}^{*},\theta_{1}^{*})$ is validated against analytical kinematics: $\|\bm{p}^{*}-\bm{p}_{\text{target}}\|_{2}\leq\epsilon\quad\text{with}\quad\bm{p}^{*}=[l_{1}^{*}\cos\theta_{1}^{*},\ l_{1}^{*}\sin\theta_{1}^{*}]^{\top}$ This provides the framework for basic robotic optimization.
IV-B Case Study 2: 2-DoF Manipulator for Workspace Optimization
@@
Discretization and quantum state encoding are performed on the combined parameter vector $\bm{z}=[\bm{q}^{\top},\bm{l}^{\top}]^{\top}=[\theta_{1},\theta_{2},l_{1},l_{2}]^{\top}$ using component-wise discretization as in (2). The full space is encoded into $N=\sum n_{i}$ qubits via uniform superposition using (3). The QML model uses a parameterized quantum circuit to approximate forward kinematics, as in Eq. (5), where the observable is position. Training minimizes the discrepancy between $\bm{p}_{k}^{\text{QML}}$ and $f_{\text{FK}}(\bm{q}^{(k)},\bm{l}^{(k)})$ across the space, as specified by $\mathcal{L}(\bm{\theta})=\frac{1}{M}\sum_{k=1}^{M}\|\bm{p}_{k}^{\text{QML}}-f_{\text{FK}}(\bm{q}^{(k)},\bm{l}^{(k)})\|^{2}$ where $k$ indexes discretized configurations, $M=2^{N}$, $\bm{p}_{k}^{\text{QML}}$ is the predicted position, $f_{\text{FK}}$ is analytical forward kinematics, and $\bm{q}^{(k)}$, $\bm{l}^{(k)}$ are decoded from $\lvert k\rangle$. The cost Hamiltonian is:
@@
with $\alpha_{R}=0$ for this planar task. Grover-optimized search amplifies states satisfying $\|f_{\text{QML}}(\lvert k\rangle)-\bm{p}_{\text{target}}\|^{2}\leq\epsilon$. The oracle is as in Eq. (12), and diffusion is $D=2\lvert\psi_{0}\rangle\langle\psi_{0}\rvert-I$, where $\lvert\psi_{0}\rangle$ is uniform superposition and II is the identity. The measured bitstring $b^{*}$ is decoded into $(\bm{q}^{*},\bm{l}^{*})$ through the inverse:
@@
IV-C Case Study 3: Dual-Arm Grasping for Coordinated Manipulation
Extending to multi-robot systems, we consider a planar dual-arm grasping task using two 2-DoF manipulators interacting with a circular object (Fig. 5(c)). The objective is to optimize the joint angles of both manipulators ($\bm{q}_{1}=[\theta_{11},\theta_{12}]^{\top}$ for the first arm and $\bm{q}_{2}=[\theta_{21},\theta_{22}]^{\top}$ for the second) to minimize the grasping error, defined as the aggregate deviation of contact points from ideal antipodal positions on the object’s surface. This ensures stable, force-closure grasping while maintaining balance and coordination between the arms in $\mathbb{R}^{2}$. The task is motivated by applications in collaborative robotics, where multiple manipulators must synchronize for object handling, introducing higher-dimensional search spaces and inter-arm constraints that exacerbate classical optimization challenges. The quantum pipeline follows the five stages adapted for the dual-arm setup: 1) Discretization & Encoding: The combined parameter vector $\bm{z}=[\bm{q}_{1}^{\top},\bm{q}_{2}^{\top}]^{\top}$ (assuming fixed link lengths for simplicity, though extensible) is discretized component-wise as in (2), encoded into $N=4\times n_{\theta}$ qubits (with $n_{\theta}$ qubits per angle), yielding $M=2^{N}$ configurations. 2) QML Kinematic Approximation: Parameterized circuit $U_{\text{QML}}(\bm{\theta})$ learns the forward kinematics for both arms, mapping to end-effector positions $\bm{p}_{1}$ and $\bm{p}_{2}$, with observables extracting contact-point data relative to the object center. 3) Cost Hamiltonian Formulation: The grasping-error observable is defined as: $\hat{H}_{\text{cost}}=\sum_{k=0}^{M-1}\left(\|\bm{p}_{1k}-\bm{c}_{\text{ideal1}}\|^{2}+\|\bm{p}_{2k}-\bm{c}_{\text{ideal2}}\|^{2}\right)\lvert k\rangle\langle k\rvert$, where $\bm{c}_{\text{ideal1}}$ and $\bm{c}_{\text{ideal2}}$ are antipodal points on the object, emphasizing coordinated minimization of deviations for stable grasp. 4) Grover-Optimized Search: Oracle $\mathcal{O}_{\text{grasp}}$ marks valid configurations: $\mathcal{O}_{\text{grasp}}\lvert k\rangle=-\lvert k\rangle\quad\text{iff}\quad\left(\|\bm{p}_{1k}^{\text{QML}}-\bm{c}_{\text{ideal1}}\|^{2}+\|\bm{p}_{2k}^{\text{QML}}-\bm{c}_{\text{ideal2}}\|^{2}\right)\leq\epsilon,$ with iterations $\sim\lfloor(\pi/4)\sqrt{M/m}\rfloor$ to amplify solutions satisfying grasp stability. 5) Classical Verification: Decoded joint angles $(\bm{q}_{1}^{*},\bm{q}_{2}^{*})$ are validated against analytical FK for both arms: $e_{\text{actual}}=\|\bm{p}_{1}^{*}-\bm{c}_{\text{ideal1}}\|^{2}+\|\bm{p}_{2}^{*}-\bm{c}_{\text{ideal2}}\|^{2}\leq\epsilon,$ with additional checks for collision avoidance between arms and force-closure conditions. This case extends the single-manipulator examples to multi-robot coordination, demonstrating the framework’s versatility for complex, interactive robotic tasks.

### v2

@@ III-B2 QML for Kinematic Approximation
where $U_{\text{QML}}(\bm{\theta})$ is a parameterized quantum circuit designed to entangle qubits and capture complex kinematic relationships—analogous to the role of neural network layers in robotic learning tasks; $\bm{p}_{k}^{\text{QML}}$ is the QML-predicted position for configuration $\lvert k\rangle$; and $\hat{\bm{P}}$ is the position observable corresponding to $\bm{p}(\bm{q},\bm{d})$ in the classical formulation. The quantum model is trained by minimizing the discrepancy with the true FK:
@@
where $f_{\text{FK}}(\bm{q}^{(k)},\bm{d}^{(k)})\|^{2}$ is the true FK. This QML optimization [31] uses gradient descent on classical computers to tune $\bm{\theta}$, while quantum hardware evaluates circuit outputs—similar to training a robotic simulator to match real-world kinematics, but leveraging quantum parallelism.
III-B3 Robotic Cost Hamiltonian Formulation
@@
Here, the Hamiltonian $\hat{H}$ acts like a potential energy landscape in robotics, where lower eigenvalues correspond to better configurations—quantum measurement samples from this landscape. See Table I for alternative observables.
The cost Hamiltonian is:
$\hat{H}_{\text{cost}}=w\cdot f_{\text{IK}}\bigl(\bm{q},\bm{d},\bm{p}_{\text{target}}\bigr),$ (8)
@@
where $\alpha_{p},\alpha_{R}>0$ balance translational and rotational units (e.g., $\alpha_{p}=1$, $\alpha_{R}=1$ empirically scaled for task-specific precision) and $\epsilon$ is the admissible pose error (set to 1-5 mm for position and 1-3° for orientation, based on typical robotic requirements). Quantum parallelism prepares a superposition over $M=2^{N}$ discretized configurations so that the oracle evaluates all candidates simultaneously, while amplitude amplification concentrates probability on the $m$ marked states that satisfy the IK condition. The number of Grover iterations is chosen near the optimum $K\approx\bigl\lfloor(\pi/4)\sqrt{M/m}\,\bigr\rfloor$, yielding quadratic speedup over classical search. For general optimization, the oracle employs the cost Hamiltonian:
@@
The parameter choices in our framework are carefully justified based on robotic requirements and quantum hardware constraints. The 9-qubit discretization provides sufficient resolution for robotic precision while remaining feasible on current NISQ devices. The error tolerance $\epsilon$ is set according to typical robotic task requirements.
III-B4 Classical Verification and Feasibility Assessment
@@
$e_{\text{actual}}=\bigl\|\,f_{\text{FK}}^{\text{analytical}}(\bm{q}^{*},\bm{d}^{*})-\bm{T}_{\text{target}}\,\bigr\|_{F}.$ (11)
Input: Pre-trained QML circuit $U_{\text{QML}}(\bm{\theta}^{*})$, target pose $\bm{T}_{\text{target}}$, error tolerance $\epsilon$
Output: Optimal configuration $(\bm{q}^{*},\bm{d}^{*})$
// Quantum State Preparation
Prepare uniform superposition over NN qubits: $\lvert\psi\rangle\leftarrow H^{\otimes N}\lvert 0\rangle^{\otimes N}$;
// Grover Search
Estimate solution count $m$ and set $K\leftarrow\left\lfloor\frac{\pi}{4}\sqrt{2^{N}/m}\right\rfloor$;
for $k\leftarrow 1$ to $K$ do
 // Oracle Application: Mark the Solution
   Apply $\mathcal{O}:\lvert x\rangle\mapsto(-1)^{f(x)}\lvert x\rangle$ where $f(x)=1$ iff $\|U_{\text{QML}}(\bm{\theta}^{*})\lvert x\rangle-\bm{p}_{\text{target}}\|^{2}\leq\epsilon$;
 // Diffusion: Amplify Marked States
   Apply $D=2\lvert\psi_{0}\rangle\langle\psi_{0}|-I$ to amplify solution probabilities;
// Measurement
Measure final state $\lvert\psi\rangle$ to obtain bitstring $b^{*}$;
Decode $b^{*}$ to classical parameters: $z_{i}^{*}=z_{i}^{\min}+\frac{k_{i}^{*}}{2^{n_{i}}-1}(z_{i}^{\max}-z_{i}^{\min})$;
Validate solution using analytical FK: $e_{\text{actual}}=\|f_{\text{FK}}^{\text{analytical}}(\bm{q}^{*},\bm{d}^{*})-\bm{T}_{\text{target}}\|_{F}$;
if $e_{\text{actual}}\leq\epsilon$ and constraints satisfied then
 return $(\bm{q}^{*},\bm{d}^{*})$
else
 return error: Adjust $\epsilon$ or retrain QML;
Algorithm 1 Quantum Optimization for IK
In the general case, the oracle marks low-cost states through the cost Hamiltonian, subsuming the IK predicate. After computing $e_{\text{actual}}$, feasibility is confirmed by checking hard constraints (joint limits, closures, collisions) and task tolerance. Failing candidates prompt adjustment of $\epsilon$, discretization, or QML training before reapplying Grover. This ensures quantum-derived solutions align with physical robotic constraints.
@@ IV Case Studies
IV-A One-DoF Manipulator for Pose Optimization
We first demonstrate the quantum optimization pipeline on a planar, single-revolute-joint manipulator. The objective is to find a link length $l_{1}$ and joint angle $\theta_{1}$ such that the end-effector position $\bm{p}=[l_{1}\cos\theta_{1},\ l_{1}\sin\theta_{1}]^{\top}$ matches a target $\bm{p}_{\text{target}}$ within a tolerance $\epsilon$. The process follows five stages:
1) Discretization & Encoding:** The continuous parameters $l_{1}\in[l_{\min},l_{\max}]$ and $\theta_{1}\in[0,2\pi)$ are discretized into $M=2^{N}$ configurations using $n_{l}$ and $n_{\theta}$ qubits, respectively ($N=n_{l}+n_{\theta}$), creating a discrete search space. 2) QML Kinematic Approximation: A parameterized quantum circuit $U_{\text{QML}}(\bm{\theta})$ is trained to approximate the forward kinematics, with an observable $\hat{\bm{P}}$ yielding the predicted position $\bm{p}^{\text{QML}}$. 3) Cost Hamiltonian Formulation: A Hamiltonian encoding the position error is constructed:
$\hat{H}_{\text{cost}}=\alpha_{p}\sum_{k=0}^{M-1}\|\bm{p}_{k}^{\text{QML}}-\bm{p}_{\text{target}}\|^{2}\lvert k\rangle\langle k\rvert$.
4) Grover-Optimized Search: Grover’s algorithm uses an oracle $\mathcal{O}_{\text{IK}}$ that marks states where $\|\bm{p}_{k}^{\text{QML}}-\bm{p}_{\text{target}}\|\leq\epsilon$, amplifying the probability of these valid solutions. 5) Classical Verification: The top solution $(l_{1}^{*},\theta_{1}^{*})$ from the quantum search is classically verified against the analytical model to ensure $\|[l_{1}^{*}\cos\theta_{1}^{*},\ l_{1}^{*}\sin\theta_{1}^{*}]^{\top}-\bm{p}_{\text{target}}\|\leq\epsilon$.
This simple case establishes a foundational and verifiable benchmark for the quantum optimization pipeline.
IV-B Two-DoF Manipulator for Workspace Optimization
@@
Discretization and quantum state encoding are performed on the combined parameter vector $\bm{z}=[\bm{q}^{\top},\bm{l}^{\top}]^{\top}=[\theta_{1},\theta_{2},l_{1},l_{2}]^{\top}$ using component-wise discretization as in (2). The full space is encoded into $N=\sum n_{i}$ qubits via uniform superposition using (3). The QML model uses a parameterized quantum circuit to approximate forward kinematics, as in (5), where the observable is position. Training minimizes the discrepancy between $\bm{p}_{k}^{\text{QML}}$ and $f_{\text{FK}}(\bm{q}^{(k)},\bm{l}^{(k)})$ across the space, as specified by $\mathcal{L}(\bm{\theta})=\frac{1}{M}\sum_{k=1}^{M}\|\bm{p}_{k}^{\text{QML}}-f_{\text{FK}}(\bm{q}^{(k)},\bm{l}^{(k)})\|^{2}$ where $k$ indexes discretized configurations, $M=2^{N}$, $\bm{p}_{k}^{\text{QML}}$ is the predicted position, $f_{\text{FK}}$ is analytical forward kinematics, and $\bm{q}^{(k)}$, $\bm{l}^{(k)}$ are decoded from $\lvert k\rangle$. The cost Hamiltonian is:
@@
with $\alpha_{R}=0$ for this planar task. Grover-optimized search amplifies states satisfying $\|f_{\text{QML}}(\lvert k\rangle)-\bm{p}_{\text{target}}\|^{2}\leq\epsilon$. The oracle is as in (10), and diffusion is $D=2\lvert\psi_{0}\rangle\langle\psi_{0}\rvert-I$, where $\lvert\psi_{0}\rangle$ is uniform superposition and $I$ is the identity. The measured bitstring $b^{*}$ is decoded into $(\bm{q}^{*},\bm{l}^{*})$ through the inverse:
@@
IV-C Dual-Arm Grasping for Coordinated Manipulation
Extending to multi-robot systems, we consider a planar dual-arm grasping task using two 2-DoF manipulators interacting with a circular object (Fig. 5(c)). The objective is to optimize the joint angles of both manipulators ($\bm{q}_{1}=[\theta_{11},\theta_{12}]^{\top}$ for the first arm and $\bm{q}_{2}=[\theta_{21},\theta_{22}]^{\top}$ for the second) to minimize the grasping error, defined as the aggregate deviation of contact points from ideal antipodal positions on the object’s surface. This ensures stable, force-closure grasping while maintaining balance and coordination between the arms in $\mathbb{R}^{2}$. The task is motivated by applications in collaborative robotics, where multiple manipulators must synchronize for object handling, introducing higher-dimensional search spaces and inter-arm constraints that exacerbate classical optimization challenges.
The quantum pipeline follows the five stages adapted for the dual-arm setup: 1) Discretization & Encoding: The combined parameter vector $\bm{z}=[\bm{q}_{1}^{\top},\bm{q}_{2}^{\top}]^{\top}$ (assuming fixed link lengths for simplicity, though extensible) is discretized component-wise as in (2), encoded into $N=4\times n_{\theta}$ qubits (with $n_{\theta}$ qubits per angle), yielding $M=2^{N}$ configurations. 2) QML Kinematic Approximation: Parameterized circuit $U_{\text{QML}}(\bm{\theta})$ learns the forward kinematics for both arms, mapping to end-effector positions $\bm{p}_{1}$ and $\bm{p}_{2}$, with observables extracting contact-point data relative to the object center. 3) Cost Hamiltonian Formulation: The grasping-error observable is defined as: $\hat{H}_{\text{cost}}=\sum_{k=0}^{M-1}\left(\|\bm{p}_{1k}-\bm{c}_{\text{ideal1}}\|^{2}+\|\bm{p}_{2k}-\bm{c}_{\text{ideal2}}\|^{2}\right)\lvert k\rangle\langle k\rvert$, where $\bm{c}_{\text{ideal1}}$ and $\bm{c}_{\text{ideal2}}$ are antipodal points on the object, emphasizing coordinated minimization of deviations for stable grasp. 4) Grover-Optimized Search: Oracle $\mathcal{O}_{\text{grasp}}$ marks valid configurations: $\mathcal{O}_{\text{grasp}}\lvert k\rangle=-\lvert k\rangle\quad\text{iff}\quad\left(\|\bm{p}_{1k}^{\text{QML}}-\bm{c}_{\text{ideal1}}\|^{2}+\|\bm{p}_{2k}^{\text{QML}}-\bm{c}_{\text{ideal2}}\|^{2}\right)\leq\epsilon,$ with iterations $\sim\lfloor(\pi/4)\sqrt{M/m}\rfloor$ to amplify solutions satisfying grasp stability. 5) Classical Verification: Decoded joint angles $(\bm{q}_{1}^{*},\bm{q}_{2}^{*})$ are validated against analytical FK for both arms: $e_{\text{actual}}=\|\bm{p}_{1}^{*}-\bm{c}_{\text{ideal1}}\|^{2}+\|\bm{p}_{2}^{*}-\bm{c}_{\text{ideal2}}\|^{2}\leq\epsilon,$ with additional checks for collision avoidance between arms and force-closure conditions. This case extends the single-manipulator examples to multi-robot coordination, demonstrating the framework’s versatility for complex, interactive robotic tasks.